OpenStudy (anonymous):

Identity help!! tan2x=

OpenStudy (anonymous):

identity means wat? @madigrieve

OpenStudy (anonymous):

a true relation?

OpenStudy (anonymous):

\[\tan(2x) = \frac{2 \tan(x)}{1 - \tan^2(x)}\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

This is the basic identity..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Or I can show you if you want..

OpenStudy (anonymous):

i get that one now

OpenStudy (anonymous):

ill post another problem hold on

OpenStudy (anonymous):

tan (A + B) = (tan A + tan B)/(1 - (tan A)(tan B)) u can write tan2 x=tan(x+x) now plug a=b=x thus tan(2x)=2tan(x)/1−tan2(x)

OpenStudy (anonymous):

Remember there is another formula: \[\tan(x +y) = \frac{\tan(x) + \tan(y)}{1 - \tan(x) \cdot \tan(y)}\] here replace y by x..

OpenStudy (anonymous):

oh ok i get the tan one now!

OpenStudy (anonymous):

that is true... \[\frac{1}{2}(\sin(x) \cos(y) + \cancel{\cos(x) \sin(y)} + \sin(x) \cos(y) - \cancel{\cos(x) \sin(y)})\] \[\implies \frac{1}{2}(2 \sin(x) \cos(y)) \implies \color{blue}{\sin(x) \cos(y)}\]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

Welcome dear..